Tuesday, November 3, 2015

GOC Service Update - Tuesday, November 10, 2015 at 14:00 UTC

The GOC will upgrade the following services beginning Tuesday, November 10, 2015 at 14:00 UTC. The GOC reserves 8 hours in the unlikely event that unexpected problems are encountered.

Some unused packages will be removed to stop MadDash from forwarding visitors to the toolkit, and the OMD/MadDash servers will be rebooted. HA mechanisms will be used to minimize end user impact.

* Ticket Blog Feed
The ticket RSS feed will change to use Oauth authentication for blog posts.

* Ticket Exchange Certificate
The Fermi certificate on tx2 will be updated to be ready for tx2 to serve as backup.